which expression is (3-2i)(-4+i)simplified?
A.-1-i
B.7-3i
C.-12+2i
D.-10+11i

Answer by stanbon(75887)

which expression is (3-2i)(-4+i)simplified?
---
= -12 +3i+8i -2(i)^2
---
= -12 + 11i +2
---
= -10 + 11i
===================
Cheers,
Stan H.
===================
A.-1-i
B.7-3i
C.-12+2i
D.-10+11i